Find x intercepts of the polynomial:

p(x)=(x-5)(x+2)(x-1)-30; . Show all steps
Answer #1

p(x) = (x-5)(x+2)(x-1) - 30

to find x intercepts we set polynomial to 0

(x-5)(x+2)(x-1) - 30 = 0

(x-5)(x+2)(x-1) = 30

x^3 - 11x^2 + 28x - 20 = 0

solving we get

x =5.801

hence, x intercepts are (5.801 , 0 )